OpenStudy (anonymous):

11.1 Vector s =(3.20 m)i -(3.70 m) j+ (2.80 m)k describes the rectilinear displacement of a body. A constant value force F =(6.70 N)i -(8.30 N)j +(4.30 N)k is applied to the body during its motion. Calculate the work done by the force during the movement.

OpenStudy (anonymous):

Work done =F.s = 3.2*6.7 + 3.7*8.3 + 2.8*4.3 = 21.44 + 30.71 + 12.04 = 64.19 Nm